# Self-bound quark stars with a first-order two-to-three flavor phase transition

### v1

@@ Abstract
We investigate self–bound quark stars in a flavor–dependent quark–mass density–dependent (QMDD) model with an excluded–volume correction. We chart the parameter space at zero pressure to identify self–bound regimes, including cases with a first–order $ud\!\to\!uds$ transition, and construct cold, $\beta$–equilibrated stellar sequences via the Tolman–Oppenheimer–Volkoff equations. The excluded–volume parameter $\kappa$ controls the stiffness of the equation of state and thus masses, radii, tidal deformabilities, and moments of inertia. Intermediate repulsion typically reconciles $M_{\max}\!\gtrsim\!2\,M_{\odot}$ with current radius/tidal constraints, with hybrid self-bound objects more compatible than pure strange–quark stars. We identify three EOS–insensitive trends— dimensionless moment of inertia vs. compactness, tidal deformability vs. compactness, and gravitational vs. baryonic compactness—whose explicit $\kappa$–dependence largely disappears in dimensionless form (the first two being notably tighter than the third). These results provide model-guided priors and tools for discriminating between hadronic and self-bound EOS families with multimessenger data.
@@ I Introduction
Strange quark stars consist of strange quark matter (SQM), a self-bound deconfined phase of up, down, and strange quarks. The concept traces back to the seminal works of Itoh [Itoh:1970uw], Bodmer [Bodmer:1971we], and Witten [Witten:1984rs], who proposed that the inclusion of the strange flavor can lower the energy per baryon below that of ordinary nuclear matter—the strange-matter hypothesis. In this picture, quark matter may constitute a stable phase with finite density at vanishing pressure. These considerations motivate a broader view of possible compact-star compositions and, in particular, prompt the question of whether self-bound phases might arise even in the absence of strangeness.
@@
Although the inclusion of the strange degree of freedom often lowers the energy per baryon—favoring a three-flavor self-bound phase—the viability of self-bound two-flavor phases cannot be dismissed a priori. This has led to studies assessing whether two-flavor quark matter (udQM)—composed of up and down quarks only—could be absolutely stable and thus self-bound under specific conditions [Holdom:2017gdc]. Even if bulk udQM satisfies $E/A<930\,\mathrm{MeV}$ at $T=P=0$, this is not, in itself, incompatible with the stability of ordinary nuclei: finite-size contributions from surface and curvature energies penalize small quark droplets. Within the multiple–reflection–expansion (MRE) framework, these finite-size corrections raise the energy per baryon by amounts that scale as $\propto\sigma A^{-1/3}$ and $\propto\gamma A^{-2/3}$, respectively, where $A$ is the baryon number and $\sigma$ ($\gamma$) denotes the surface (curvature) tension [Madsen:1998uh]. As a result, nuclei can remain absolutely stable even if bulk udQM is the ground state: the decay of a nucleus into a finite $ud$-quark drop must overcome the surface and curvature barriers. Quantitatively, these effects imply a minimum baryon number $A_{\min}$ above which $ud$-quark matter droplets are favored, while hadronic matter is preferred below it; representative estimates yield $A_{\min}\!\gtrsim\!300$ [Holdom:2017gdc]. The precise threshold depends on the microphysics that controls $\sigma$ and $\gamma$, for which effective-model calculations typically find sizable values, reinforcing the stabilizing role of finite-size terms [Lugones:2013ema, Lugones:2020qll]. These considerations parallel the classical strange-matter arguments [Witten:1984rs, Farhi:1984qu] while emphasizing that stability in bulk does not automatically extend to finite lumps. In the stellar context, macroscopic self-bound configurations can circumvent finite-size penalties and may be realized as compact stars composed of deconfined quark matter, whereas in vacuum small $ud$ droplets are unstable and hadronize.
@@
Ab initio QCD at finite baryon density is presently out of reach: lattice calculations suffer from the sign problem, and perturbation theory is trustworthy only at asymptotically large densities. We therefore employ phenomenological equations of state (EOS). The MIT bag model encodes confinement through a bag constant $B$ but lacks dynamical chiral symmetry breaking/restoration, whereas the Nambu–Jona-Lasinio (NJL) model captures chiral dynamics via local four-fermion interactions yet does not confine [MITmodificado1, MITmodificado2, LugonesInterativo, NJL1, NJL2, Buballa:1998pr]. This complementarity motivates a framework that incorporates, in a minimal and consistent way, both confinement and (effective) chiral behavior. In this work we adopt the quark-mass density–dependent (QMDD) model, wherein quark effective masses acquire a density dependence. Enforcing thermodynamic consistency introduces additional terms in the pressure that act as a confining contribution, while the masses approach their current values at high density, mimicking chiral restoration [Lugones:2022upj, Lugones:2023zfd, Lugones:2024ryz].
@@
Two variants of the QMDD scheme were introduced in Ref. [Lugones:2022upj]: a flavor-blind ansatz, where the effective mass depends on the baryon number density, and a flavor-dependent ansatz, where each flavor’s mass tracks its own particle number density. A comprehensive survey of the parameter space shows that, in the flavor-blind case, among parameter sets that yield absolute stability, the vanishing-pressure ground state is invariably three-flavor. In contrast, the flavor-dependent formulation admits a broad domain of parametrizations in which two-flavor matter is self-bound at zero pressure.
Most of our previous astrophysical applications employed the flavor-blind variant [Lugones:2023zfd, Lugones:2024ryz, Lugones2025universality]. Here we focus on the flavor-dependent scheme. For parameter sets whose zero-pressure ground state is two-flavor, our preliminary calculations robustly produce a first-order $ud\!\to\!uds$ transition at higher pressures. This transition reshapes the EOS and permits compact configurations with an udsuds core enveloped by a $ud$ mantle—a quark–quark “hybrid” that we term self-bound hybrid stars. In what follows we analyze the flavor-dependent model in detail and quantify the astrophysical consequences of the $ud\!\to\!uds$ transition.
@@
The paper is organized as follows. Section II introduces the flavor-dependent QMDD framework, detailing the density-dependent mass ansatz and the excluded-volume prescription, and summarizes the resulting zero-temperature EOS. In Section III we delineate the stability window in the $(a,C)$ plane and discuss causality bounds. Section IV presents the thermodynamics and phase structure—Gibbs energy per baryon, $p(\epsilon)$, and speed of sound—quantifying the two- to three-flavor transition (transition pressure and energy-density jump). In Section V we compute global stellar properties by solving the Tolman–Oppenheimer–Volkoff (TOV) equations and confront mass–radius, tidal deformability, and moment-of-inertia predictions with current multimessenger constraints. Section VI discusses EOS-insensitive (“universal”) relations for self-bound stars. We conclude by summarizing the main results and their implications in Sec. VII. Technical details of the EOS construction and of the computations of Love numbers and the moment of inertia are collected in the Appendices A and B.
@@ II The flavor-dependent QMDD model
We model cold quark matter as a gas of non-interacting quasiparticles of flavors $u,d,s$ with effective masses $M_{i}$, supplemented by free electrons to ensure charge neutrality. Within the QMDD framework [Lugones:2022upj], interaction effects are encoded through density-dependent quark masses while the expressions for the energy density and the number densities retain the free-gas functional form. Specifically, the effective mass of each flavor is taken as
@@
where $i\in\{u,d,s\}$, $m_{i}$ is the (constant) current mass of flavor $i$, $n_{i}$ is its number density, and $C$ and $a$ are flavor-independent parameters. Units are chosen so that $M_{i}$ is in MeV when $n_{i}$ is expressed in $\mathrm{fm}^{-3}$; throughout we adopt natural units with $\hbar=c=1$. In contrast to our earlier “flavor-blind” formulation—where $M_{i}$ depended only on the baryon number density—we employ here a flavor-dependent mass function $M_{i}(n_{i})$.
The limiting behavior of Eq. (1) reproduces the expected qualitative features: as $n_{i}\!\to\!0$, $M_{i}$ diverges, suppressing quasiparticle excitations and thereby mimicking confinement; at asymptotically large densities, $M_{i}\!\to\!m_{i}$, providing a phenomenological realization of (effective) chiral restoration, particularly if $m_{i}=0$. The exponent $a$ can be related to a schematic interquark potential, $v(r)\propto r^{a}$, which controls how rapidly medium effects diminish with increasing density [Lugones:2022upj].
A natural way to encode repulsive interactions in the flavor-dependent QMDD model is to introduce a phenomenological excluded-volume correction that penalizes arbitrarily high local number densities, thereby stiffening the EOS at moderate and high densities [Lugones:2023zfd]. Consistent with the flavor-dependent ansatz in Eq. (1), we take quarks of flavor $i$ to reduce only the volume available to the same flavor (no interspecies reduction):
@@
We set $b(n_{i})$ via an ansatz that reflects the expected microphysical behavior of repulsive interactions: it yields a substantial per-particle excluded volume at low to intermediate densities and vanishes at high densities, consistent with asymptotic freedom. For consistency with the flavor-dependent mass prescription in Eq. (1), we also take the excluded volume per particle to be flavor dependent. Accordingly, in analogy with Refs. [Lugones:2023zfd, Lugones:2022upj], we adopt
@@
Substituting Eqs. (1) and (4) into the general flavor-dependent QMDD expressions of Ref. [Lugones:2023zfd] yields the complete EOS. We collect the resulting formulas in Appendix A.
@@ III Stability window
Figure 1 shows the stability window of the flavor-dependent QMDD model in the $(a,C)$ plane. The curves $e_{ud}^{0}=930\,\mathrm{MeV}$ (black) and $e_{uds}^{0}=930\,\mathrm{MeV}$ (red) delimit where two- and three-flavor matter, respectively, are self-bound at zero pressure: points lying below the former (latter) yield self-bound udud (udsuds) matter. Accordingly, self-bound $ud$ quark matter (udQM) is realized when $e_{ud}^{0}<e_{uds}^{0}$, whereas SQM occurs for parameter choices with $e_{uds}^{0}<e_{ud}^{0}$. Points above both 930 MeV curves correspond to hybrid EOSs—hadronic at low pressure and deconfined only beyond a transition pressure—leading to stars with a quark core and a hadronic envelope. An additional curve, $c_{s}^{0}=c$, marks parameter choices for which the sound speed at zero pressure equals the speed of light. To its left, the EOS remains causal down to $p=0$; to its right, the $p\!\to\!0$ limit becomes acausal ($c_{s}^{0}>c$), thereby excluding self-bound quark matter. Although the flavor-dependent QMDD model involves three parameters—$a$, $C$, and the excluded-volume parameter bb—Ref. [Lugones:2023zfd] shows that $b$ does not modify the zero-pressure energy per baryon and thus does not affect the self-bound criterion. The stability region is therefore fully determined by $(a,C)$.
@@
We analyze four benchmark points in the $(a,C)$ plane, indicated by (a)–(d) in Fig. 1 and listed in Table 1. Point (a), $(a,C)=(5.5,40)$, lies in the SQM sector and serves as the reference strange–quark–matter EOS. Points (b) $(1,135)$, (c) $(2,70)$, and (d) $(3,60)$ sample the self-bound $ud$ domain and exemplify self-bound hybrid configurations in which a two– to three–flavor transition can occur. Unless stated otherwise, each benchmark is evaluated at several excluded–volume strengths, typically $\kappa=0,\,0.3,\,0.6$ [Lugones:2023zfd]; the chosen $\kappa$ values will be specified in the corresponding figures and captions.

### v2

@@ Abstract
We investigate self–bound quark stars in a flavor–dependent quark–mass density–dependent model with an excluded–volume correction. We chart the parameter space at zero pressure to identify self–bound regimes, including parametrizations in which self–bound two–flavor matter undergoes a genuine first–order $ud\to uds$ transition at finite pressure. We construct cold, $\beta$–equilibrated stellar sequences and compute the corresponding global properties (mass–radius relation, tidal deformability, and moment of inertia). For a wide region of the model parameter space, we find that the onset of a udsuds core occurs before the maximum–mass configuration is reached, yielding self–bound hybrid stars that follow the typical strange–quark–star sequence morphology but develop a characteristic kink at $p_{c}=p_{\rm tr}$ along the stellar curves. The excluded–volume parameter $\kappa$ controls the stiffness of the equation of state and thus masses, radii, tidal deformabilities, and moments of inertia; intermediate repulsion typically reconciles $M_{\max}\!\gtrsim\!2\,M_{\odot}$ with current astrophysical constraints. We further identify two equation–of–state–insensitive trends—dimensionless moment of inertia versus compactness and gravitational versus baryonic compactness. These results provide model–guided priors and tools for discriminating between hadronic and self–bound equations of state with multimessenger data.
@@ I Introduction
Strange quark stars consist of strange quark matter (SQM), a self-bound deconfined phase of up, down, and strange quarks. The concept traces back to the seminal works of Itoh [11], Bodmer [2], and Witten [32], who proposed that the inclusion of the strange flavor can lower the energy per baryon below that of ordinary nuclear matter—the strange-matter hypothesis. In this picture, quark matter may constitute a stable phase with finite density at vanishing pressure. These considerations motivate a broader view of possible compact-star compositions and, in particular, prompt the question of whether self-bound phases might arise even in the absence of strangeness.
@@
Although the inclusion of the strange degree of freedom often lowers the energy per baryon—favoring a three-flavor self-bound phase—the viability of self-bound two-flavor phases cannot be dismissed a priori. This has led to studies assessing whether two-flavor quark matter (udQM)—composed of up and down quarks only—could be absolutely stable and thus self-bound under specific conditions [9]. Even if bulk udQM satisfies $E/A<930\,\mathrm{MeV}$ at $T=P=0$, this is not, in itself, incompatible with the stability of ordinary nuclei: finite-size contributions from surface and curvature energies penalize small quark droplets. Within the multiple–reflection–expansion (MRE) framework, these finite-size corrections raise the energy per baryon by amounts that scale as $\propto\sigma A^{-1/3}$ and $\propto\gamma A^{-2/3}$, respectively, where $A$ is the baryon number and $\sigma$ ($\gamma$) denotes the surface (curvature) tension [22]. As a result, nuclei can remain absolutely stable even if bulk udQM is the ground state: the decay of a nucleus into a finite $ud$-quark drop must overcome the surface and curvature barriers. Quantitatively, these effects imply a minimum baryon number $A_{\min}$ above which $ud$-quark matter droplets are favored, while hadronic matter is preferred below it; representative estimates yield $A_{\min}\!\gtrsim\!300$ [9]. The precise threshold depends on the microphysics that controls $\sigma$ and $\gamma$, for which effective-model calculations typically find sizable values, reinforcing the stabilizing role of finite-size terms [14, 15]. These considerations parallel the classical strange-matter arguments [32, 7] while emphasizing that stability in bulk does not automatically extend to finite lumps. In the stellar context, macroscopic self-bound configurations can circumvent finite-size penalties and may be realized as compact stars composed of deconfined quark matter, whereas in vacuum small $ud$ droplets are unstable and hadronize.
@@
Ab initio QCD at finite baryon density is presently out of reach: lattice calculations suffer from the sign problem, and perturbation theory is trustworthy only at asymptotically large densities. We therefore employ phenomenological equations of state (EOS). The MIT bag model encodes confinement through a bag constant $B$ but lacks dynamical chiral symmetry breaking/restoration, whereas the Nambu–Jona-Lasinio (NJL) model captures chiral dynamics via local four-fermion interactions yet does not confine [13, 12, 16, 25, 26, 4]. This complementarity motivates a framework that incorporates, in a minimal and consistent way, both confinement and (effective) chiral behavior. In this work we adopt the quark-mass density–dependent (QMDD) model, wherein quark effective masses acquire a density dependence. Enforcing thermodynamic consistency introduces additional terms in the pressure that act as a confining contribution, while the masses approach their current values at high density, mimicking chiral restoration [17, 18, 20].
@@
Two variants of the QMDD scheme were introduced in Ref. [17]: a flavor-blind ansatz, where the effective mass depends on the baryon number density, and a flavor-dependent ansatz, where each flavor’s mass tracks its own particle number density. A comprehensive survey of the parameter space shows that, in the flavor-blind case, among parameter sets that yield absolute stability, the vanishing-pressure ground state is invariably three-flavor. In contrast, the flavor-dependent formulation admits a broad domain of parametrizations in which two-flavor matter is self-bound at zero pressure [17].
In the flavor-dependent scheme, parameter sets whose zero-pressure ground state is two-flavor necessarily become three-flavor at sufficiently high pressure, since opening a new Fermi sea eventually renders the strange sector energetically favored. This provides a natural motivation to analyze how the $ud\!\to\!uds$ conversion proceeds and what sets its characteristic scales. In the flavor-dependent scheme, the dynamical masses can vary more independently than in the flavor-blind case, because each flavor responds to its own number density rather than to a common baryon-density scale. As a result, strangeness can be strongly suppressed at low densities while $u$ and $d$ quarks are not suppressed to the same extent, allowing self-bound two-flavor matter at $p\!\to\!0$ and delaying the onset of $s$ quarks to higher pressure. In fact, the exploratory study of Ref. [18] indicates that the appearance of $s$ quarks need not be gradual: it can occur abruptly through a genuine first-order phase transition. This motivates a more systematic investigation of the $ud\!\to\!uds$ scenario and its astrophysical implications.
Most of our previous astrophysical applications employed the flavor-blind variant [18, 20, 19]. Here we focus on the flavor-dependent scheme and analyze, across the relevant stability window, the stellar sequences implied by parametrizations whose zero-pressure ground state is two-flavor. We solve for cold, $\beta$-equilibrated configurations and quantify how the ensuing quark–quark transition impacts global observables, in particular the mass–radius relation, tidal deformability, and moment of inertia, for the resulting self-bound hybrid stars. Unlike conventional hybrid stars based on a hadron–to–quark phase transition, the stars considered here are self–bound at all pressures and contain only quark degrees of freedom. At the stellar level, the phase transition gives rise to configurations consisting of a udsuds core surrounded by a $ud$ mantle. This structure is qualitatively different from that obtained in models with a smooth crossover between quark phases, where no density discontinuity is present. The resulting self–bound hybrid stars occupy a distinct region of the mass–radius plane and exhibit characteristic behavior in their tidal deformabilities and related observables, which we analyze in detail below.
Very recently, hybrid quark stars featuring quark–quark phase transitions have also been investigated in Refs. [33, 34]. In particular, Yang et al. [33] considered compact stars containing a $ud\to uds$ interface, while Zhang et al. [34] explored related slow–stable (in the sense of Ref. [21]) self–bound hybrid configurations within different effective descriptions. Although these studies address a broadly similar physical scenario, their modeling strategy differs substantially from ours. Both works [33, 34] rely on highly parameterized, piecewise prescriptions for the EOS, in which the transition is implemented between two independently modeled phases. As a result, key characteristics—such as the transition pressure and the associated energy–density discontinuity—are not derived from a single, unified microphysical framework, and the origin of the $ud\to uds$ threshold is not traced back to an underlying dynamical mechanism. By contrast, in our framework the $ud\to uds$ transition emerges dynamically from the particle–number–density dependence of the quark masses, without introducing a separate phase–transition construction by hand. Moreover, we quantify how excluded–volume effects reshape the stellar sequences and identify equation–of–state–insensitive universal relations specific to self–bound hybrid stars. These aspects allow us to complement and extend the recent literature by highlighting both the microphysical origin of the transition and its distinct astrophysical implications.
@@
The paper is organized as follows. Section II introduces the flavor-dependent QMDD framework, detailing the density-dependent mass ansatz and the excluded-volume prescription, and summarizes the resulting zero-temperature EOS. In Section III we delineate the stability window in the $(a,C)$ plane and discuss causality bounds. Section IV presents the thermodynamics and phase structure—Gibbs energy per baryon, $p(\epsilon)$, and speed of sound—quantifying the two- to three-flavor transition (transition pressure and energy-density jump). In Section V we construct cold, $\beta$-equilibrated stellar sequences by integrating the Tolman–Oppenheimer–Volkoff equations and the standard perturbation equations for tidal deformability and slow rotation (moment of inertia), and we compare the resulting mass–radius relations with current multimessenger constraints. Section VI discusses EOS-insensitive (“universal”) relations for self-bound stars. We conclude by summarizing the main results and their implications in Sec. VII. Technical details of the EOS construction and of the computations of Love numbers and the moment of inertia are collected in the Appendices A and B.
@@ II The flavor-dependent QMDD model
We model cold quark matter as a gas of non-interacting quasiparticles of flavors $u,d,s$ with effective masses $M_{i}$, supplemented by free electrons to ensure charge neutrality. Within the QMDD framework [17], interaction effects are encoded through density-dependent quark masses while the expressions for the energy density and the number densities retain the free-gas functional form. Specifically, the effective mass of each flavor is taken as
@@
where $i\in\{u,d,s\}$, $m_{i}$ is the (constant) current mass of flavor $i$, $n_{i}$ is its number density, and $C$ and $a$ are considered flavor-independent parameters.
We note that the flavor-dependent QMDD scheme already embodies $SU(3)_{f}$ breaking at the dynamical level through the strange-quark current mass and the resulting distinct density evolution of the $s$ sector under $\beta$ equilibrium and charge neutrality. In this sense, the model is “flavor-dependent” because each quasiparticle mass tracks the number density of its own flavor, $M_{i}=M_{i}(n_{i})$, rather than a common baryon-density scale as in flavor-blind variants. This reduced correlation among the density dependences of $M_{u}$, $M_{d}$, and $M_{s}$ plays a central role in enabling domains where strangeness is suppressed at low pressure while $u$ and $d$ quarks are not, thereby opening the possibility that $ud$ matter is self-bound at $p\!\to\!0$ and that the onset of $s$ quarks is delayed to higher pressures.
At the same time, Eq. (1) adopts a minimal parametrization in which the mass-functional parameters $(a,C)$ are taken to be flavor-independent. Allowing explicit parameter flavor dependence, e.g., $C_{s}\neq C_{u,d}$ and/or $a_{s}\neq a_{u,d}$, would further enlarge the model space and may shift quantitative outcomes such as the transition density/pressure and the size of the udsuds core. We do not pursue that extension here in order to keep the parameter survey tractable and to isolate the physical consequences of the flavor-dependent density evolution itself; explicit flavor dependence in $(a,C)$ constitutes a natural next step beyond the scope of the present work (see also the exploratory study in the Appendix of Ref. [17]).
Units are chosen so that $M_{i}$ is in MeV when $n_{i}$ is expressed in $\mathrm{fm}^{-3}$; throughout we adopt natural units with $\hbar=c=1$. In contrast to our earlier “flavor-blind” formulation—where $M_{i}$ depended only on the baryon number density—we employ here a flavor-dependent mass function $M_{i}(n_{i})$.
The limiting behavior of Eq. (1) reproduces the expected qualitative features: as $n_{i}\!\to\!0$, $M_{i}$ diverges, suppressing quasiparticle excitations and thereby mimicking confinement; at asymptotically large densities, $M_{i}\!\to\!m_{i}$, providing a phenomenological realization of (effective) chiral restoration, particularly if $m_{i}=0$. The exponent $a$ can be related to a schematic interquark potential, $v(r)\propto r^{a}$, which controls how rapidly medium effects diminish with increasing density [17].
A natural way to encode repulsive interactions in the flavor-dependent QMDD model is to introduce a phenomenological excluded-volume (EV) correction that penalizes arbitrarily high local number densities, thereby stiffening the EOS at moderate and high densities [18]. In the absence of an underlying Lagrangian with explicit mediator fields, the EV prescription provides an economical effective parametrization of short-range repulsive channels that are expected to be operative in dense, nonperturbative QCD matter. Such repulsion is also indicated empirically by the existence of compact stars with masses $\gtrsim 2\,M_{\odot}$, which generically requires a sufficiently stiff high-density EOS. In this sense, EV plays here a role analogous to vector interactions in field-theoretical descriptions (e.g., relativistic mean-field models for hadronic matter or NJL/vector-MIT extensions for quark matter), while remaining consistent with the defining structure of QMDD models based on density-dependent mass functionals.
Consistent with the flavor-dependent ansatz in Eq. (1), we take quarks of flavor $i$ to reduce only the volume available to the same flavor (no interspecies reduction):
@@
We set $b(n_{i})$ via an ansatz that captures the expected qualitative behavior of an effective repulsion: it suppresses excessively large local densities and increases the pressure at a given energy density, thereby stiffening the EOS at moderate densities, while becoming negligible at very high densities, in line with asymptotic freedom. In the present implementation, the repulsive correction is incorporated through the available-volume factor in Eq. (2), i.e., through a finite-volume (“hard-core”) effect akin to Van der Waals–type treatments. For consistency with the flavor-dependent mass prescription in Eq. (1), we also allow the excluded volume per particle to depend on the density of the same flavor. Accordingly, in analogy with Refs. [18, 17], we adopt
@@
Substituting Eqs. (1) and (4) into the general flavor-dependent QMDD expressions of Ref. [18] yields the complete EOS, in which the EV correction effectively mimics repulsive interactions without introducing additional mediator fields and couplings. We collect the resulting formulas in Appendix A.
@@ III Stability window
Figure 1 shows the stability window of the flavor-dependent QMDD model in the $(a,C)$ plane. The curves $e_{ud}^{0}=930\,\mathrm{MeV}$ (black) and $e_{uds}^{0}=930\,\mathrm{MeV}$ (red) delimit where two- and three-flavor matter, respectively, are self-bound at zero pressure: points lying below the former (latter) yield self-bound udud (udsuds) matter. Accordingly, self-bound $ud$ quark matter (udQM) is realized when $e_{ud}^{0}<e_{uds}^{0}$, whereas SQM occurs for parameter choices with $e_{uds}^{0}<e_{ud}^{0}$. Points above both 930 MeV curves correspond to hybrid EOSs—hadronic at low pressure and deconfined only beyond a transition pressure—leading to stars with a quark core and a hadronic envelope. An additional curve, $c_{s}^{0}=c$, marks parameter choices for which the sound speed at zero pressure equals the speed of light. To its left, the EOS remains causal down to $p=0$; to its right, the $p\!\to\!0$ limit becomes acausal ($c_{s}^{0}>c$), thereby excluding self-bound quark matter. Because excluded-volume repulsion stiffens the EOS at finite density, one might expect it to also affect the $c_{s}^{0}=c$ boundary. In our implementation, however, the $p\!\to\!0$ sound speed is independent of the excluded-volume (cf. Ref. [18]): we have verified numerically that $c_{s}^{0}$ is invariant under changes in $\kappa$, so the causality line in Fig. 1 is determined solely by $(a,C)$. This behavior is consistent with the structure of the EV prescription: the quark-sector thermodynamics is obtained from the pointlike expressions by a constant available-volume factor and a density rescaling, which leaves the ratio $dp/d\epsilon$ unchanged in the $p\!\to\!0$ limit. Consequently, varying $\kappa$ does not expand the acausal region and does not alter whether a given benchmark point is excluded by the $c_{s}^{0}>c$ criterion. Therefore, although the flavor-dependent QMDD model involves three parameters—$a$, $C$, and the excluded-volume parameter $b$, the stability region is fully determined by $(a,C)$.
@@
We analyze four benchmark points in the $(a,C)$ plane, indicated by (a)–(d) in Fig. 1 and listed in Table 1. Point (a), $(a,C)=(5.5,40)$, lies in the SQM sector and serves as the reference strange–quark–matter EOS. Points (b) $(1,135)$, (c) $(2,70)$, and (d) $(3,60)$ sample the self-bound $ud$ domain and exemplify self-bound hybrid configurations in which a two– to three–flavor transition can occur. Unless stated otherwise, each benchmark is evaluated at several excluded–volume strengths, typically $\kappa=0,\,0.3,\,0.6$ [18]; the chosen $\kappa$ values will be specified in the corresponding figures and captions.